OpenStudy (anonymous):

Use dimensional analysis to complete each equation. 18 in. = ■ ft

sammixboo (sammixboo):

Well there is 12 inches in a foot, correct?

OpenStudy (xapproachesinfinity):

12 inch =1 feet so 1 inch =1/12 feet

OpenStudy (anonymous):

hw can i get my answr

OpenStudy (kainui):

You can multiply anything by 1 and it doesn't change it, so you take this equation here: \[\Large 12 in = 1 ft\] then you divide both sides by 12 in to get: \[\Large 1 = \frac{1 ft}{12 in}\] That lets you multiply for example \[\Large 24 in * 1 = 24 in * \frac{1 ft}{12 in} = 2 ft\]

OpenStudy (anonymous):

so what my answer